# Model Poisoning Attacks to Federated Learning via Multi-Round Consistency

### v2

@@ Abstract
Model poisoning attacks are critical security threats to Federated Learning (FL). Existing model poisoning attacks suffer from two key limitations: 1) they achieve suboptimal effectiveness when defenses are deployed, and/or 2) they require knowledge of the model updates or local training data on genuine clients. In this work, we make a key observation that their suboptimal effectiveness arises from only leveraging model-update consistency among malicious clients within individual training rounds, making the attack effect self-cancel across training rounds. In light of this observation, we propose PoisonedFL, which enforces multi-round consistency among the malicious clients’ model updates while not requiring any knowledge about the genuine clients. Our empirical evaluation on five benchmark datasets shows that PoisonedFL breaks eight state-of-the-art defenses and outperforms seven existing model poisoning attacks. Moreover, we also explore new defenses that are tailored to PoisonedFL, but our results show that we can still adapt PoisonedFL to break them. Our study shows that FL systems are considerably less robust than previously thought, underlining the urgency for the development of new defense mechanisms.
@@ 1 Introduction
Existing model poisoning attacks suffer from two key limitations. First, they [8, 10, 11, 9] only leverage in-round consistency, i.e., they send the same or similar malicious model updates in a single training round to amplify the attack effect, but neglect multi-round consistency. For instance, Fang attack [10] crafts the malicious model updates in a given training round to maximally deviate the global model in that round, while MPAF [9] sets the malicious model updates in a training round as the magnified difference between the current global model and a random model. Such a strategy limits the attack effectiveness when defenses are deployed. This is because defenses weaken the attack effect in every training round by filtering and clipping the malicious model updates; and the weakened attack effects further self-cancel across training rounds due to the lack of multi-round consistency, as shown by our measurement results in Section 4. Second, most existing attacks [8, 10, 11] require knowledge of genuine clients’ model updates or local training data. They often assume compromising a significant number of genuine clients to obtain such knowledge, which is notably impractical. These attacks with suboptimal effectiveness and/or impractical assumptions give a false sense of FL security.
@@
Our work:  In this work, we propose PoisonedFL, a novel model poisoning attack to FL, which overcomes the above limitations of existing attacks. Specifically, to enhance attack effectiveness, PoisonedFL leverages two key components: 1) multi-round consistency, and 2) dynamic attack magnitude adjustment. The first component ensures that even if the attack effect is weakened in individual training rounds by defenses, the cumulative attack effect over multiple training rounds remains undeterred. The second component aims to prevent the malicious model updates from being entirely filtered out in individual training rounds by defenses. Moreover, PoisonedFL does not require genuine clients’ information, and is agnostic to the defenses deployed by the server.
@@
For the first component, our key observation is that a model, when updated substantially in a random direction, exhibits significantly degraded accuracy (see Fig. 10 in Appendix). Based on this observation, we formulate an all-round optimization problem that facilitates multi-round consistency. Specifically, the objective of our optimization problem is to maximize the magnitude of the total aggregated model updates in a random update direction, which is characterized by a random sign vector $\bm{s}$. Each dimension of the sign vector $\bm{s}$ is either +1 or -1 uniformly at random, meaning that the total aggregated model updates increase or decrease the corresponding parameter of the global model. With such total aggregated model updates, the final global model is substantially moved along the random update direction, leading to a large testing error rate. Note that our objective stands substantially distinct from those of existing attacks, as it enforces multi-round consistency without self-cancellation. However, it is challenging to solve the all-round optimization problem in a given round because the model updates in future training rounds are not available. To address the challenge, we transform the optimization problem as a per-round one, which aims to find malicious model updates on the malicious clients to maximize the magnitude of the aggregated model update along the update direction $\bm{s}$ in each round. In particular, our malicious model update is a dimension-wise product of the sign vector $\bm{s}$ and a magnitude vector, a dimension of which indicates the magnitude of the corresponding dimension of the malicious model update.
@@
For the second component, we propose to dynamically adjust the magnitude vector in each training round based on the attack results in the past rounds, making our attack effective against the unknown defense that the server may have deployed. Our key idea is to find a magnitude vector that has a large magnitude while avoiding the malicious model update being filtered out by the unknown defense. On one hand, based on the global models and malicious model updates in the previous training rounds, we estimate the model updates of genuine clients to determine an unit magnitude vector. The magnitude vector is the product of the unit magnitude vector and a scaling factor. On the other hand, we set the scaling factor based on whether the attack achieves the objective in the past rounds. In particular, our attack objective is that the total aggregated model updates have the random update direction $\bm{s}$. We leverage statistical hypothesis testing to check whether our attack in the previous rounds achieves this objective. If not, it means the malicious updates were filtered out by the defense in previous rounds and thus we decrease the scaling factor to be more stealthy.
We evaluate our PoisonedFL on five benchmark datasets. Our results show that, under the least knowledge scenario – that is, no knowledge about the defense nor genuine local training data/models – PoisonedFL can break eight state-of-the-art FL defenses [12, 13, 14, 15, 16, 17, 18]. For instance, PoisonedFL can reduce the global models to random guessing in many cases. Moreover, in most cases, PoisonedFL achieves higher testing error rates than existing attacks [8, 10, 11, 9] when a defense is deployed even if we assume the attacker has access to the local models of all genuine clients. In the remaining cases, both PoisonedFL and some existing attacks reduce the global models to random guessing. Furthermore, PoisonedFL remains effective across different FL settings including the fraction of malicious clients, the degree of non-IID of the genuine clients’ local training data, the number of epochs used to train a local model, and the fraction of clients selected to participate in each training round.
We also explore new defenses tailored to PoisonedFL. For instance, PoisonedFL enforces the malicious model updates across training rounds to have the same sign vector, but model updates from genuine clients may not satisfy such constraints. Therefore, we design a method to detect fake clients based on the sign vectors of the model updates across multiple training rounds. However, we can still adapt PoisonedFL to counter such defense with a minor impact on attack effectiveness. In particular, instead of using the same sign vector across all training rounds, our adapted PoisonedFL randomly flips a small portion of the sign vector in each training round when crafting the malicious model updates. Our results highlight the need for new defense mechanisms.
@@
  • We uncover a fundamental limitation of existing model poisoning attacks, i.e., their inconsistent malicious model updates across multiple training rounds self-cancel the attack effect, leading to suboptimal attack effectiveness.
@@ 2.1 Federated Learning (FL)
Federated Learning (FL) [1, 2] is a machine learning paradigm that allows many clients (e.g., smartphones) to collaboratively train a model without sending their raw local training data to a server. Suppose we have $n$ clients and each client has a local training dataset $\mathcal{D}_{i}$, $i=1,2,\cdots,n$. FL aims to solve an optimization problem that is defined as $\min\sum_{i=1}^{n}\mathcal{L}{(\bm{w},\mathcal{D}_{i})}$, where $\bm{w}\in\mathbb{R}^{d}$ is the model (called global model), $d$ is the number of parameters in the model $\bm{w}$, and $\mathcal{L}{(\bm{w},\mathcal{D}_{i})}$ is the local loss function for client $i$. FL iteratively trains the global model $\bm{w}$. In the $t$-th training round, the server first distributes the current global model $\bm{w}^{t-1}$ to all clients or a subset of them. For simplicity, we assume all clients are selected, but we will also evaluate a subset of them are selected in each training round in our experiments. Each client $i$ then trains a local model $\bm{w}_{i}^{t}$ based on the current global model and its local training data. For instance, client $i$ initializes $\bm{w}_{i}^{t}$ as $\bm{w}^{t-1}$ and iteratively updates $\bm{w}_{i}^{t}$ via minimizing $\mathcal{L}{(\bm{w},\mathcal{D}_{i})}$ using Stochastic Gradient Descent (SGD). After training the local model $\bm{w}_{i}^{t}$, client $i$ sends the model update $\bm{g}^{t}_{i}=\bm{w}^{t}_{i}-\bm{w}^{t-1}$ to the server. Finally, the server aggregates the received model updates and updates the global model as follows:
@@
where $\bm{g}^{t}={{AR}}(\{\bm{g}^{t}_{i}\}_{i\in[1,n]})$ is the aggregated model update, ${AR}$ denotes the aggregation rule used by the server, and $\{\bm{g}^{t}_{i}\}_{i\in[1,n]}$ is the set of clients’ model updates. For instance, when ${AR}$ is FedAvg [2], the aggregated model update ${{AR}}(\{\bm{g}^{t}_{i}\}_{i\in[1,n]})$ is the average of the clients’ model updates.
@@ 2.2 Poisoning Attacks to FL
In model poisoning attacks [19, 8, 10, 11], malicious clients corrupt the global model via sending carefully crafted malicious model updates to the server. In a targeted attack [19], a poisoned global model only misclassifies attacker-chosen inputs with certain properties (e.g., embedded with an attacker-chosen trigger). In contrast, in an untargeted attack [8, 10, 11], a poisoned global model indiscriminately misclassifies a large number of clean inputs, leading to a large testing error rate and eventually denial-of-service attacks. Therefore, we focus on untargeted attacks in this work.
Requiring genuine clients’ information:  One category of model poisoning attacks [8, 10, 11] require that an attacker has access to genuine clients’ model updates or local training data. To obtain such knowledge, these attacks often assume an attacker can compromise a large fraction of genuine clients. For instance, in a little is enough (LIE) attack [8], the attacker computes the average of the genuine model updates of the compromised genuine clients and adds perturbation to them to craft the malicious model updates in each training round. Concurrent to LIE, Fang et al. [10] proposed the first optimization-based attack framework that can be tailored to any aggregation rule. Their key idea is that the attacker crafts the malicious model updates such that the aggregated model update after attack deviates substantially from the before-attack one in each training round. Building upon this attack framework [10], the authors in [11] assume the malicious model update to be the sum of the before-attack aggregated model update and a fixed perturbation vector scaled by a factor, where the scaling factor is optimized to maximize the deviation between the after-attack aggregated model update and the before-attack one.
Not requiring genuine clients’ information:  The other category of model poisoning attacks [9] do not rely on any genuine clients’ information. Therefore, they only assume that an attacker can inject fake clients into an FL system, which is more realistic than compromising genuine clients. Since genuine clients’ information is not available, these attacks craft malicious model updates based on heuristics. For instance, in MPAF [9], the attacker crafts malicious model updates such that the global model progressively moves towards a random model in each training round.
Limitations of existing attacks:  Existing attacks suffer from the following limitations. First, the first category of attacks require an attacker to compromise a large fraction of genuine clients, which is not affordable in production FL [20]. Second, both categories of attacks craft malicious model updates independently in different training rounds, leading to self-cancellation of attack effect across training rounds. As a result, they achieve suboptimal attack effectiveness when defenses are deployed.
@@ 2.3 Defenses against Poisoning Attacks
Byzantine-robust aggregation rules:  Instead of using the average of the clients’ model updates to update the global model, Byzantine-robust aggregation rules [12, 13, 21, 15, 16, 17, 22, 23] typically employ statistical analysis to filter out or clip outlier model updates before aggregating them. For instance, Multi-Krum [12] iteratively selects the model update that is the closest to its nearest neighbor model updates, and aggregates them. Dimension-wise Median (or Trimmed Mean) [17] uses the median (or trimmed-mean) of the clients’ local models for each dimension. Norm Bound [16] aims to minimize the impact of potential malicious model updates by clipping the magnitudes of all model updates. FLTrust [13] assumes the server uses a clean and small dataset to bootstrap trust. Specifically, in FLTrust [13], the server trains a server model and assigns trust scores to clients by measuring the similarity between the server model update and clients’ model updates. Under some assumptions about the learning problem and clients’ local training data distribution, these Byzantine-robust aggregation rules often can theoretically guarantee that the learnt global model does not change much when the number of malicious clients is bounded. However, they are still vulnerable to our attack because these assumptions often do not hold and a bounded change in the parameters of the global model does not guarantee a small change of error rate.
Provably robust defenses:  Provably robust defenses [24, 14, 25] can provide stronger theoretical guarantees than Byzantine-robust aggregation rules. In particular, given a number of malicious clients, a provably robust defense can guarantee a lower bound of testing accuracy no matter what poisoning attacks the malicious clients use. For instance, FLCert [14] trains multiple global models, each of which is trained using a subset of clients. Given a testing input, FLCert predicts its label by taking a majority vote among the multiple global models. A major limitation of these provably robust defenses is that they can only tolerate a small number of malicious clients, i.e., the lower bound of testing accuracy reduces to 0 when the number of malicious clients is moderate. As a result, they are still vulnerable to our attack since an attacker can easily inject a large number of fake clients.
Malicious clients detection:  These defenses [26, 18] aim to detect malicious clients during or after the training process. After detecting malicious clients, the server can further efficiently recover/re-train an accurate global model using the remaining clients [27]. For instance, FLDetector [18] aims to detect malicious clients by utilizing the inconsistency of a client’s model updates across multiple training rounds. Their key idea is that the model updates of a benign client in multiple training rounds are all calculated based on the client’s local training data, while the model updates of a malicious client are crafted without following the learning procedure. As a result, the model updates of a benign client satisfy certain consistency constraints across training rounds, while those of a malicious client do not, making it possible to distinguish between benign and malicious clients. In our attack, we enforce the malicious model updates of a malicious client to be consistent across training rounds, making FLDetector ineffective.
@@ 3 Threat Model
Attacker’s goal:  Following previous works  [8, 10, 11], we consider the goal of an attacker is to poison the training process of FL such that the final learnt global model makes incorrect predictions for a large number of clean inputs indiscriminately. In other words, the poisoned global model has a large testing error rate. Such a global model would eventually be abandoned, leading to denial-of-service attacks [28].
@@
Attacker’s capability:  We assume that the attacker is capable of injecting fake clients into an FL system as malicious clients. We note that this is a practical threat model. Multiple platforms [29, 30, 31] can be used to automatically inject fake devices/clients in a distributed system. For instance, when the FL system learns a global model on Android smartphones (i.e., clients), e.g., Google’s next-word prediction on Android devices [32], an attacker can use a laptop to simulate fake smartphones via Android emulators [9].
@@
Attacker’s background knowledge:  We consider the minimal background knowledge scenario, which represents the most realistic threat model. Specifically, the attacker has access to the global models during the training process since the server sends them to the clients including the fake ones. However, the attacker does not have access to genuine local training data nor local models on any genuine client. Moreover, we assume the attacker does not know the defense (e.g., Byzantine-robust aggregation rule, provably robust defense, or detector) deployed by the server.
We note that our threat model is the same as previous fake client based model poisoning attacks [9], which do not require genuine clients’ information.
4 Motivation
(a) Testing error rate
(b) Flipping rate
We first introduce our key observation about why existing model poisoning attacks achieve suboptimal attack effectiveness, which motivates the design of our PoisonedFL in the next section. We show experimental results on the MNIST dataset with the default parameter settings described in Section 6.1. We assume the server uses Trimmed Mean [17] as the defense. Moreover, we assume the ratio of fake clients to genuine clients is $20\%$. We consider MPAF [9], an attack not requiring genuine clients’ information, and Fang [10], a representative attack that requires genuine clients’ information. To give advantages to Fang, we assume the attacker knows the local models of all participating genuine clients in each round when crafting malicious model updates on fake clients.
In each training round, these attacks craft a malicious model update for each fake client selected by the server to participate in training. A malicious model update $\bm{g}_{i}^{t}$ of fake client $i$ in training round $t$ can be decomposed into a dimension-wise product of a sign vector $\bm{s}_{i}^{t}$ and a magnitude vector $|\bm{g}_{i}^{t}|$, where ||$\cdot$|| ⋅ | means dimension-wise absolute value of a vector. Each dimension of $\bm{s}_{i}^{t}$ is either +1 or -1, while all dimensions of $|\bm{g}_{i}^{t}|$ are non-negative. A dimension with a +1 (or -1) sign aims to increase (or decrease) the corresponding dimension/parameter of the global model. For a fake client, we say a dimension is flipped in round $t$ if its sign in the malicious model update is flipped, compared to that in the previous training round. A flipped dimension in a training round $t$ may cancel the attack effect of the previous training round $t-1$. This is because the malicious model updates aim to increase the dimension of the global model in one training round but decrease it in the other. For each training round $t$, we define flipping rate as the fraction of flipped dimensions of a malicious model update averaged over the fake clients.
Fig. 1(a) shows the testing error rate of the global model as a function of training round when no attack, MPAF, or Fang is used; while Fig. 1(b) shows the flipping rate as a function of training round when MPAF or Fang is used to craft malicious model updates on the fake clients. We observe that the flipping rate of MPAF is always larger than 35% and that of Fang is larger than 40% in the last around 1,500 training rounds, leading to self-cancellation of attack effect in many rounds. For instance, in Fang attack, the flipping rate is around 10% in nearly the first 1,000 training rounds, i.e., the malicious model updates maintain some degree of “consistency”. As a result, the global model under Fang attack has a large testing error rate in those rounds as shown in Fig. 1(a). However, the flipping rate rapidly increases as the training proceeds, eventually leading to self-cancellation of attack effect and an accurate global model.

### v3

@@ Abstract
Model poisoning attacks are critical security threats to Federated Learning (FL). Existing model poisoning attacks suffer from two key limitations: 1) they achieve suboptimal effectiveness when defenses are deployed, and/or 2) they require knowledge of the model updates or local training data on genuine clients. In this work, we make a key observation that their suboptimal effectiveness arises from only leveraging model-update consistency among malicious clients within individual training rounds, making the attack effect self-cancel across training rounds. In light of this observation, we propose PoisonedFL, which enforces multi-round consistency among the malicious clients’ model updates while not requiring any knowledge about the genuine clients. Our empirical evaluation on five benchmark datasets shows that PoisonedFL breaks eight state-of-the-art defenses and outperforms seven existing model poisoning attacks. Our study shows that FL systems are considerably less robust than previously thought, underlining the urgency for the development of new defense mechanisms. Our source code is available at https://github.com/xyq7/PoisonedFL/.
@@ 1 Introduction
Existing model poisoning attacks suffer from two key limitations. First, they [10, 17, 32, 13] only leverage in-round consistency, i.e., they send the same or similar malicious model updates in a single training round to amplify the attack effect, but neglect multi-round consistency. Fig. 1 reveals that many parameters’ attack directions (+1+1/1-1) reverse from the previous round (a large flipping rate), leading to self-cancellation and diminished impact after extensive training (small testing error rate). Detailed setups and discussion on this observation is deferred to Appendix A. Second, most existing attacks [10, 17, 32] require knowledge of genuine clients’ model updates or local training data. They often assume compromising a significant number of genuine clients to obtain such knowledge, which is notably impractical. These attacks with suboptimal effectiveness and/or impractical assumptions give a false sense of FL security.
@@
Our work:  In this work, we propose PoisonedFL, a novel model poisoning attack to FL, which overcomes the above limitations of existing attacks. Specifically, to enhance attack effectiveness, PoisonedFL leverages two key components: 1) multi-round consistency, and 2) dynamic attack magnitude adjustment. The first component ensures that even if the attack effect is weakened in individual training rounds by defenses, the cumulative attack effect over multiple training rounds remains undeterred. The second component aims to prevent the malicious model updates from being entirely filtered out in individual training rounds by defenses. Moreover, PoisonedFL does not require genuine clients’ information, and is agnostic to the defenses deployed by the server, thus can be performed on injected fake clients with least information of the FL system.
@@
For the first component, our key observation is that a model, when updated substantially in a random direction, exhibits significantly degraded accuracy (see Fig. 6 in Appendix). Based on this observation, we formulate an optimization problem that maximizes the magnitude of the total aggregated model updates in a random update direction, which is characterized by a random and fixed sign vector $\bm{s}$. With such total aggregated model updates, the final global model is substantially moved along the random update direction, leading to a large testing error rate. To achieve this, we propose a simple attack strategy in each round, where the malicious model update is a dimension-wise product of the fixed sign vector $\bm{s}$ and a dynamic non-negative magnitude vector, a dimension of which indicates the magnitude of the corresponding dimension of the malicious model update. Note that our objective stands substantially distinct from those of existing attacks, as it enforces multi-round consistency without self-cancellation.
@@
For the second component, we propose to dynamically adjust the magnitude vector to have a large magnitude while avoiding the malicious model update being filtered out by the unknown defense. On one hand, based on the global models and malicious model updates in the previous training rounds, we estimate the model updates of genuine clients to determine an unit magnitude vector. The magnitude vector is the product of the unit magnitude vector and a scaling factor. On the other hand, we set the scaling factor based on whether the attack achieves the objective in the past rounds. In particular, our attack objective is that the total aggregated model updates have a random update direction $\bm{s}$. We leverage statistical hypothesis testing to check whether our attack in the previous rounds achieves this objective. If not, it means the malicious model updates were filtered out by the defense in previous rounds and thus we decrease the scaling factor to be more stealthy.
We evaluate our PoisonedFL on five benchmark datasets. Our results show that, under the least knowledge scenario – that is, no knowledge about the defense nor genuine local training data/models – PoisonedFL can break eight state-of-the-art FL defenses [11, 14, 16, 28, 34, 39, 40] in different FL settings. We also explore new defenses tailored to PoisonedFL. However, we can still adapt PoisonedFL to counter such defense with a minor impact on attack effectiveness. Our results highlight the need for new defense mechanisms.
@@
  • We uncover a fundamental limitation of existing model poisoning attacks, i.e., their inconsistent malicious model updates across multiple rounds self-cancel the attack effect, leading to suboptimal effectiveness.
@@ 2.1 Federated Learning (FL)
Suppose we have $n$ clients and each client has a local training dataset $\mathcal{D}_{i}$, $i=1,2,\cdots,n$. FL iteratively trains the global model $\bm{w}$. In the $t$-th training round, the server first distributes the current global model $\bm{w}^{t-1}$ to all clients or a subset of them. For simplicity, we assume all clients are selected, but we will also evaluate a subset of them are selected in each training round in our experiments. Each client $i$ then trains a local model $\bm{w}_{i}^{t}$ based on the current global model and its local training data. For instance, client $i$ initializes $\bm{w}_{i}^{t}$ as $\bm{w}^{t-1}$ and iteratively updates $\bm{w}_{i}^{t}$ via minimizing $\mathcal{L}{(\bm{w},\mathcal{D}_{i})}$ using Stochastic Gradient Descent (SGD). After training the local model $\bm{w}_{i}^{t}$, client $i$ sends the model update $\bm{g}^{t}_{i}=\bm{w}^{t}_{i}-\bm{w}^{t-1}$ to the server. Finally, the server aggregates the received model updates and updates the global model as follows:
@@
where $\bm{g}^{t}={{AR}}(\{\bm{g}^{t}_{i}\}_{i\in[1,n]})$ is the aggregated model update, AR{AR} denotes the aggregation rule used by the server, and $\{\bm{g}^{t}_{i}\}_{i\in[1,n]}$ is the set of clients’ model updates. For instance, when AR{AR} is FedAvg [26], the aggregated model update $\bm{g}^{t}$ is the average of the clients’ model updates.
@@ 2.2 Poisoning Attacks to FL
In model poisoning attacks [8, 10, 17, 32], malicious clients corrupt the global model via sending carefully crafted malicious model updates to the server. This work focuses on untargeted attacks [10, 17, 32], where a poisoned global model indiscriminately misclassifies a large number of clean inputs, leading to a large testing error rate and eventually denial-of-service attacks.
Requiring genuine clients’ information:  One category of model poisoning attacks [10, 17, 32] requires that an attacker has access to genuine clients’ model updates or local training data. To obtain such knowledge, these attacks often assume an attacker can compromise a large fraction of genuine clients. For instance, a little is enough (LIE) attack [10] manipulates malicious model updates with small perturbation to the genuine model updates, while Fang et al. [17] propose to craft malicious model updates towards the reverse direction of genuine model updates.
Not requiring genuine clients’ information:  The other category of model poisoning attacks [13] do not rely on any genuine clients’ information. Therefore, they only assume that an attacker can inject fake clients into an FL system, which is more realistic than compromising genuine clients. For instance, in MPAF [13], the attacker crafts malicious model updates on fake clients towards the initial model.
Limitations of existing attacks:  Existing attacks suffer from the following limitations. First, the first category of attacks requires an attacker to compromise a large fraction of genuine clients, which is not affordable in production FL [33]. Second, both categories of attacks suffer from self-cancellation of attack effect across training rounds. As a result, they achieve suboptimal attack effectiveness when defenses are deployed.
@@ 2.3 Defenses against Poisoning Attacks
Byzantine-robust aggregation rules:  Instead of using the average of the clients’ model updates to update the global model, Byzantine-robust aggregation rules [11, 14, 27, 28, 34, 39, 18, 19] typically employ statistical analysis to filter out or clip outlier model updates before aggregating them. For instance, Multi-Krum [11] selects model update with vector-wise analysis, while Median (or Trimmed Mean) [39] uses dimension-wise median (or trimmed-mean) of the clients’ local models. Norm Bound [34] clips the magnitudes of all model updates. FLTrust [14] assumes the server uses a clean and small dataset to bootstrap trust. Under some assumptions about the learning problem and clients’ local training data distribution, these Byzantine-robust aggregation rules often can theoretically guarantee that the learnt global model does not change much when the number of malicious clients is bounded. However, they are still vulnerable to our attack because these assumptions often do not hold and a bounded change in the parameters of the global model does not guarantee a small change of error rate.
Provably robust defenses:  Provably robust defenses [15, 16, 36] provide theoretical guarantees of a lower bound of testing accuracy given a number of malicious clients. For instance, FLCert [16] trains multiple global models using a subset of clients and predicts with a majority vote. A major limitation of these provably robust defenses is that they can only tolerate a small number of malicious clients, i.e., the lower bound of testing accuracy reduces to 0 when the number of malicious clients grows (e.g., 20%). As a result, they are still vulnerable to our attack since an attacker can easily inject some fake clients.
Malicious clients detection:  These defenses [25, 40, 37] aim to detect malicious clients during or after the training process. For instance, FLDetector [40] detects malicious clients by utilizing the inconsistency of a client’s model updates across multiple training rounds. However, in our attack, we enforce the malicious model updates of a malicious client to be consistent across training rounds, making FLDetector ineffective.
@@ 3 Threat Model
Attacker’s goal:  Following previous works  [10, 17, 32], we consider the goal is to poison the training process of FL such that the final learnt global model makes incorrect predictions for clean inputs indiscriminately. In other words, the poisoned global model has a large testing error rate. Such a global model would eventually be abandoned, leading to denial-of-service attacks [9].
@@
Attacker’s capability:  We assume that the attacker can inject fake clients into an FL system and control them as malicious clients. We note that this is a practical threat model. Multiple platforms [5, 1, 2] can be used to automatically inject fake devices/clients in a distributed system. For instance, when the FL system learns a global model on Android smartphones (clients), e.g., Google’s next-word prediction on Android devices [3], an attacker can use a laptop to simulate fake smartphones via Android emulators [13].
@@
Attacker’s background knowledge:  We consider the minimal background knowledge scenario, which represents the most realistic threat model. Specifically, the attacker has access to the global models during the training since the server sends them to all clients. However, the attacker has no access to genuine local training data nor local models on any genuine client. Moreover, we assume the attacker does not know the defense deployed by the server. We note that our threat model is the same as previous fake client based model poisoning attacks [13].